OpenStudy (anonymous):

Find the square root of -25i.

OpenStudy (anonymous):

A number.

OpenStudy (anonymous):

It could be interpreted as an expression; you may have thought i was a variable.

OpenStudy (anonymous):

For this, Let: \[-25i = x + iy\] Now tell me what is x here and what is y here by comparing..

OpenStudy (anonymous):

x is the integer, y is the complex part of of number to be multiplied by i. I have i = a^2 +2iab - b^2 and a^2 - b^2 = 0.

OpenStudy (anonymous):

After that I am stuck.

OpenStudy (anonymous):

What you get as 2ab??

OpenStudy (anonymous):

Squaring both sides.

OpenStudy (anonymous):

I am using a+bi. I suppose I should have posted it in x+yi form.

OpenStudy (anonymous):

Let: \[\sqrt{-25i} = a + ib\] Squaring both the sides: \[-25i = (a^2-b^2) + 2abi\] From here: \((a^2 - b^2) = 0\) -----------1 \[2ab = -25\] Square it: \[4a^2.b^2 = 625\] -----------2 Got till here??

OpenStudy (anonymous):

I see. So a = 25/2

OpenStudy (anonymous):

No a is not that much.. Use this formula: \[(a^2 + b^2)^2 = (a^2-b^2)^2 + 4a^2b^2\] Can you find \(a^2 + b^2\) from here??

OpenStudy (anonymous):

Where did that come from?

OpenStudy (anonymous):

Oh wait, is a = 5/sqrt(2) ?

OpenStudy (anonymous):

Yes you find a correct.. Now find b..

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

\[(a^2 - b^2)^2 = a^4 + b^4 - 2a^2.b^2\] Now add \(4a^2.b^2\) both the sides, \[(a^2-b^2)^2 + 4a^2.b^2 = a^2 + b^4 + 2a^2.b^2 \implies (a^2+b^2)^2\] This is what I have written there..

OpenStudy (anonymous):

I found b to be (5i)/sqrt(2).

OpenStudy (anonymous):

This one is wrong.. Show me how you did it??

OpenStudy (anonymous):

I mean b = 5/sqrt(2). Of a = 5/sqrt(2), then that minus itself gives 0.

OpenStudy (anonymous):

Yes now you are right.. Both the values of a and b will be same.. But there is one little mistake: \[a = \pm \frac{5}{\sqrt{2}} \qquad Or \qquad b = \pm \frac{5}{\sqrt{2}}\]

OpenStudy (anonymous):

I see. Thank you. That setting the complex number to a+bi is a clever thing.

OpenStudy (anonymous):

Um, @waterineyes you know it simplifies easier than that, right? I'm using a different technique than what you're doing, mind posting what you got?

OpenStudy (anonymous):

In simplified form I mean, I want to see if it matches what I got

OpenStudy (anonymous):

(-1)^something style

OpenStudy (anonymous):

After that you will do like this: \[\sqrt{-25i} = \pm (\frac{5}{\sqrt{2}} + i \frac{5}{\sqrt{2}})\]

OpenStudy (anonymous):

\[\sqrt{-25i} = \sqrt{-1*5*5*\sqrt{-1}} = 5 * \sqrt{-1 \sqrt{-1}}=5(-1(-1)^{1/2})^{1/2}=-5 (-1)^{3/4}\]

OpenStudy (anonymous):

Is this the square root of this??

OpenStudy (anonymous):

You have done correctly there is no doubt that this is wrong but I have learned to take the square root of complex number as I did.. Sorry..

OpenStudy (anonymous):

Yep, I'm correct. http://www.wolframalpha.com/input/?i=square+root+of+%28-25i%29

OpenStudy (anonymous):

Wolfram is not everything.. There are certain algorithms are made to solve particular things and you must follow that..

OpenStudy (anonymous):

I just used fractional exponents and treated -1 as if it was a variable for a moment, that's how I did that in my head.

OpenStudy (anonymous):

Think about it, what's \(\large\sqrt{\sqrt{x}}\) ? x\(^{1/4}\), agreed? What's 1-(1/4)? 3/4. Where did that 1 go? Well, kind of like this: (\(\sqrt{-1}\))^2 = \(((-1)^{2/4})^{4/2}\) = -1 I am following what you've said waterineyes, it's just unusual to me because I did it with a completely different method.

OpenStudy (anonymous):

That's why I asked what you got with your method :-) There's often more than one correct way of solving something in mathematics after all